A researcher knows that the average height of Filipino women is 1.53 m. A random sample of




26 women was taken and found to have a mean height if 1.56 m with a standard deviation of




0.1 m. Is there a reason to believe that the 26 women in the sample are significantly taller than




the others?

1
Expert's answer
2022-06-10T14:00:10-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le1.53"

"H_1:\mu>1.53"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=25" and the critical value for aright-tailed test is "t_c =1.708141."

The rejection region for thisright-tailed test is "R = \{t:t>1.708141\}."

The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{1.56-1.53}{0.1/\sqrt{26}}=1.5297"


Since it is observed that "t=1.5297<1.708141=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, "df=25" degrees of freedom, "t=1.5297" is "p=0.069324," and since "p=0.069324>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 1.53, at the "\alpha = 0.05" significance level.